If k x velocity x mass is displacement, determine the dimensions of “k”
1
Expert's answer
2020-10-29T07:01:15-0400

Solution

According to given theory of question


"k\times velocity\times mass" =displacement

Putting dimension of velocity and mass

"k=\frac{displacement}{velocity\times Mass}"


"k=\frac{meter}{\frac{meter\times Kg}{sec}}"


"k=\frac{sec}{Kg}" or "[M^{-1}L^{0}T^{1}]"